Solve Z(p^2+q^2)+px+qy=0
1
Expert's answer
2020-09-27T18:10:44-0400

Let f=z("p^2+q^2)+px+qy"


Using charpit's method,

"\frac{dz}{-p\frac{df}{dp}-q\frac{df}{dq}}" ="\frac{dp}{\frac{df}{dx}+p\frac{df}{dz}}" ="\frac{dq}{\frac{df}{dy}+q\frac{df}{dz}}=\frac{dx}{\frac{-df}{dp}}=\frac{dy}{\frac{-df}{dq}}"


From second and third factor we have

"\frac{dp}{\frac{df}{dx}+p\frac{df}{dz}}=\frac{dq}{\frac{df}{dy}+q\frac{df}{dz}}"

"\frac{dp}{p+p(p^2+q^2)}=\frac{dq}{q+q(p^2+q^2)}"


"\frac{dp}{p}=\frac{dq}{q}"


So by integrating ,we get

"log(p)=log(q)-log(a)"


q=pa

Putting the value of p in given equation,

"p^2(1+a^2)z+px+apy=0"


"p(1+a^2)z+(x+ay)=0"


so p="-\frac{x+ay}{z(1+a^2)}" so q="-\frac{a(x+ay)}{(1+a^2)z}"


now we have to solve

dz=pdx+qdy

"dz=-\frac{x+ay}{(1+a^2)z}dx-\frac{a(x+ay)}{1+a^2)z}dy"


"(1+a^2)zdz=-xdx-a(ydx+xdy)-a^2ydy"

On integrating the above equation, we get


"(1+a^2)\frac{z^2}{2}=-\frac{x^2}{2}-axy-a^2\frac{y^2}{2}"


"(1+a^2)z=-x^2-2axy-a^2y^2"

This is the solution.